OpenStudy (johnnydicamillo):

Find the inverse of g given g(x) = x -3 /9x+4. Find domain and rage using interval notation

OpenStudy (anonymous):

switch x and y , then solve for y

OpenStudy (johnnydicamillo):

okay @sourwing, I get this then \[x=\frac{ y-3 }{ 9y+4 }\] correct?

OpenStudy (johnnydicamillo):

should I multiply x by 9y + 4?

OpenStudy (johnnydicamillo):

I am stuck

OpenStudy (anonymous):

yes

OpenStudy (johnnydicamillo):

so will that look like \[9yx+4x = y -3\] or \[9y+4(x) = y-3\]?

OpenStudy (anonymous):

9yx + 4x = y - 3 is correct move all the y terms to one side and the remaining terms to the other side, then solve for y

OpenStudy (johnnydicamillo):

so\[8y=\frac{ -3-4x }{ x }\]?

OpenStudy (anonymous):

9yx + 4x = y - 3 9yx - y = -4x - 3 y (9x - 1) = -4x - 3 y = -(4x - 3) / (9x - 1)

OpenStudy (johnnydicamillo):

I would distribute that negative so it is \[\frac{ -4x=3 }{ 9x-1 }\]?

OpenStudy (johnnydicamillo):

I mean -4x +3